OpenStudy (anonymous):

Write the equation 0.3x2 + 5x - 7 = 0 in general form and then choose the value of "b."

OpenStudy (anonymous):

In general form the coefficient must be on whole number. Given 0.3x2 + 5x - 7 = 0 times 10 throughout the whole equation 3x^2 + 50x - 70 = 0 the b value here is 50.

OpenStudy (anonymous):

general form is ax^2+bx+c=0, and multiplying the given equation by 10 ,we get 3x^2 + 50x - 70 = 0 and hence b=50